OpenStudy (anonymous):

simplify: 1/(3-6t) + 1/(3-6t)^2

sam (.sam.):

-(2(3t-2))/(9(2t-1)^(2))

OpenStudy (anonymous):

(3-6t+1)/(3-6t)^2 =(4-6t)/(9-36t+36t^2)

OpenStudy (anonymous):

both of u above have different answers!

OpenStudy (anonymous):

if u read attentively, they are same..

OpenStudy (anonymous):

yes they are sorry

OpenStudy (anonymous):

Thanks to u both

sam (.sam.):

Use my one, mine is more simplified

OpenStudy (anonymous):

i have doubt if the expression u wrote is simplified, @sam. i guess u have to expand every term and write down their co-efficients to 'simplify'

OpenStudy (anonymous):

@sam @arnab09 : can u kindly tell me if we can split this : (3-6t)+1 / (3-6t)(3-6t) into....... (3-6t)/(3-6t) + 1/(3-6t) ??

OpenStudy (anonymous):

which becomes 1+ 1/(3-6t) ??

sam (.sam.):

(3-6t)+1 / (3-6t)(3-6t) = 1/(3-6t) + 1/(3-6t)(3-6t)

OpenStudy (anonymous):